A two-stage rocket moves in space at a constant velocity of +4070 m/s. The two stages are then separated by a small explosive charge placed between them. Immediately after the explosion the velocity of the 1200-kg upper stage is +5960 m/s. What is the velocity (magnitude and direction) of the 2030-kg lower stage immediately after the explosion?
Answer #1

by the conservation of mass we have

m1 + m2 = m

1200 + 2030 = m

m = 3230 kg

by the conservation of momentum we have

3230*4070 = 1200*5960 + 2030*v

v = 2952.75 m/s ( in the same direction as initial velocity )